how can i write this (4/5)^x+1 *(5/4)^x in scientific notation

Answer by stanbon(75887)

how can i write this (4/5)^x+1 *(5/4)^x in scientific notation
-------
---------------------------------
(4/5)^(x+1) * (5/4)^x
-----
= (4/5)^x * (4/5) * (5/4)^x
-----
= [(4/5)(5/4)]^x * (4/5)
-------------
= [1]^x*(4/5)
--------
= 1*4/5 = 4/5 = 0.8
---------
= 8.0*10^-1
==================
Cheers,
Stan H.
==================